OpenStudy (anonymous):

Write the prime factorization for 180. can someone explain how to do this?

OpenStudy (anonymous):

well, primefactorization is when you break a number down into the prime numbers that, when multiplied equal the number you started with

OpenStudy (anonymous):

do you know how to start?

OpenStudy (anonymous):

nope a teacher explained to me but it wasn't to clear for me

OpenStudy (anonymous):

ok, i like to see if the number is divisable by 2. is 180 divisable by 2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

perfect, so 180 factors to 2 * 90. 2 is a prime number so nuw you factor 90....again, is 90 divisable by 2?

OpenStudy (anonymous):

Just factor out prime numbers until you're left with a prime number at the end. 180 = 60 * 3 60 = 3 * 20 20 = 4 * 5 4 = 2 * 2 So, 180 = 3 * 3 * 5 * 2 * 2 = 3^2 * 2^2 * 5